1.The sum of three consecutive odd integers is -3. What are the integers.
2. The perimeter of a rectangle is 68 meters. The width is one-hald meter less than one half of the length. What is the area of the rectangle.

Answer by mananth(12270)

The sum of three consecutive odd integers is -3. What are the integers.
2. The perimeter of a rectangle is 68 meters. The width is one-hald meter less than one half of the length. What is the area of the rectangle.

let the integer x
next two odd integer will be x-2 and x-4
x+x-2 +x-4=-3
3x-6=-3
3x=3
x=1
2. The perimeter of a rectangle is 68 meters. The width is one-half meter less than one half of the length. What is the area of the rectangle.
let the length be x
width = x/2 -1/2
2x+ 2(x/2-1/2) =68
2x+x-1=68
3x=69
x=23
y=11